OpenStudy (anonymous):

Solve the equation. –6 + 3x = –9

OpenStudy (anonymous):

Well the goal is to separate x from the rest of the numbers. We can start by taking the -6 and moving it to the other side of the equation with +6. Do you follow?

OpenStudy (anonymous):

–6 + 3x = –9 3x=-9+6 3x=-3 x=-3/3 x=-1